OpenStudy (anonymous):

form a polynomial f(x) with real coefficient having the given degree and zeros: degree:5; zeros: 2, -3i and 4-i

myininaya (myininaya):

f(x)=(x-2)(x+3i)(x-(4-i))(x-3i)(x+(4-i))

OpenStudy (anonymous):

f(x)=(x-2)(x+3i)(x-4+i)(x-3i)(x+4-i)

myininaya (myininaya):

yep

OpenStudy (anonymous):

the answer is f(x)=x^4+2ix^3-6ix^2-18i-6x^3+17x^2-49x+72 ... is it right??

OpenStudy (anonymous):

^ why would the answer be degree 4?